# Protected gap closing and reopening in topological-insulator Josephson junctions

### v1

@@ 4 Locality of the gap closing in perpendicular magnetic fields
Next, we study the behaviour of our device in higher magnetic fields and compare the LDOS probed at TJ1 and TJ2. When a perpendicular magnetic field $B$ is applied to the junction, $\varphi$ becomes dependent on the position along the junction ($y$ direction). If $B$ is strong enough to nucleate Josephson vortices, the position where the local phase difference becomes $\varphi=(2n+1)\pi$ is identified as the Josephson vortex-core, and the gapless 1D mode in the Fu-Kane theory for global $\varphi=(2n+1)\pi$ turns into a Majorana zero-mode confined to the Josephson vortex-core[26]. The concept of this Majorana zero-mode generation is sketched in Fig. 3a. A spatially varying phase $\varphi$ along a TI Josephson junction caused by $B$ leads to a position-dependent gap function $m\propto\cos\frac{\varphi(y)}{2}$ as indicated by the colour code in the junction[26]. Associated with the sign change in the gap function, the gap closes locally at $\varphi(y)=(2n+1)\pi$ (white area), corresponding to the core of the Josephson vortex, and a Majorana zero-mode (green dots) nucleates that is confined along the junction by the phase gradient. We note that the topologically-protected nature of the gap closing is intuitively understandable in this situation: the region of the junction with $-\pi<\varphi(y)<\pi$ (red) has the even-parity ground state with $\mathbb{Z}_{2}$ index of 0, whereas the region with $\pi<\varphi(y)<3\pi$ (blue) has the odd-parity ground state with $\mathbb{Z}_{2}$ index of 1; therefore, the boundary between the two regions must become gapless for the topological invariant to change.
@@
We indeed observed the locality of this gap closing. The top and bottom panels of Fig. 3c show the $dI/dV$ data simultaneously measured at TJ1 and TJ2 as a function of $V_{b}$ and $B$ in the 5 – 6 mT range. Both junctions present a comparable evolution of their bound state spectrum, but there is a clear phase shift and their periods are slightly different. For TJ2 we find a periodicity of $200\,\mathrm{\mu T}$, smaller than that of TJ1 ($230\,\mathrm{\mu T}$). This difference in the period is consistent with the sample geometry (see supplement for details) in which the effective junction area considering the flux focusing effect[34, 35, 36] is about $\sim$10of the loop area. To understand the LDOS at TJ1 and TJ2, we have calculated the expected texture of the gap function based on the device geometry. In our junctions, the self-inductance is negligible and its width is much shorter than the Josephson penetration depth (see supplement), justifying the description of the phase gradient in terms of
@@
with $A$ the loop area, $\Phi_{0}=\frac{h}{2e}$ the magnetic flux quantum, and $L_{\rm eff}$ the effective length of the junction taking the flux-focusing effect into account. The evolution of the gap function $m\propto\cos\frac{\varphi(y)}{2}$ is calculated from this formula with the loop and junction area as fitting parameters (see supplement) and visualized in the middle panel of Fig. 3c. We show snapshots of the phase-texture in the junction for a range of magnetic fields corresponding to the experimental situation. In this range, there are two to three Josephson vortices inside the junction, whose position is tuned by the global perpendicular magnetic field. They enter the junction from the bottom where TJ2 is located, and move upwards to TJ1 with increasing magnetic field. Whenever a Josephson vortex-core crosses the location of a tunnel junction, a gap closing is observed. This demonstrates the local nature of the gap closing in the presence of Josephson vortices. The data shown in Fig. 3c was taken at an applied back-gate voltage of $V_{\rm G}=-5$ V, and we chose the magnetic field range of 5 – 6 mT for the visibility of the phase gradient. The same pattern was found to continue into higher magnetic fields until the Al superconductivity was suppressed at 15 mT (see supplement).
@@
We note that the spectra shown in Fig. 3c essentially look the same as those in Fig. 1e. This is because the Majorana zero-mode in the Josephson vortex-core is not well-separated in energy. For the conditions realized, it is accompanied by a quasi-continuum of low-energy bound states[26]. Taking the smearing due to the finite-temperature into account, the local spectrum becomes a broad peak, indistinguishable from the Fu-Kane situation. In Fig. 3b, we show the magnitude of the gap function near a Josephson vortex-core and the finite-energy bound states calculated from the model in Ref. 26, along with the spatial extent of the expected Majorana zero-mode. For the parameters relevant to our device with $\mu$ = 25 meV, the calculated finite-energy bound states form a quasi-continuum with a minigap of only 2% of $\Delta$ (i.e. 3 $\mu$eV), which is of the order of the thermal energy at the base temperature.
5 Stability of the gap closing
@@
In the next step, we check the robustness of the observed features with respect to the chemical potential, as tuned by a global back gate. Figure 4a shows the $dI/dV$ data in $B$ = 0 T at a high bias voltage ($V_{\rm b}$ = 1 mV) as a function of the back-gate voltage $V_{G}$ for both TJ1 and TJ2. The overall tendency is that the conductance decreases for more negative gate-voltages, indicating that the TI surface is in the $n$-type regime. Superimposed to this general trend, we observe broad peaks (typical peak width of a few V), which are commonly found in dielectric-based tunnelling devices and conventionally explained by resonant states in the dielectric[31]. Figure 4b presents the $V_{\rm G}$-dependencies of the normalized conductance spectra obtained at TJ1: the upper and lower panels show the evolutions of the maximum excitation gap for $\varphi=0$ (upper) and the spectrum observed for $\varphi=\pi$ (lower) in a wide range of $V_{\rm G}$ from $-20$ to $+20$ V. The maximum excitation gap is very stable, indicating that the resonance states due to impurities are not affecting the proximity-induced gap. On the other hand, the position of the conductance peak for $\varphi=\pi$ is slightly shifted to the negative $V_{\rm b}$ side at $V_{\rm G}\approx$ 12 – 20 V, which is probably due to an influence of the strong resonance states located around 13 V (see Fig. 4a). Remarkably, the broad ZBCP that accompanies the gap closing is stable for a wide range of $V_{\rm G}$ from $-10$ V to $+12$ V. It is worth noting that the ZBCP observed in this $V_{\rm G}$ range is not correlated with the weaker resonance states seen in Fig. 4a, which further supports the intrinsic origin of the gap closing. For $V_{\rm G}\lesssim-10$ V, the gap closing is not accompanied by a clear ZBCP. We believe that this is because our non-tunable tunnel contacts did not couple sufficiently well to the bound states in this regime, preventing us from resolving a peak.
@@
In Figs. 4c and 4d, we show the $B$-dependencies of the normalized conductance spectra for equally-spaced $V_{\rm G}$ values between $-20$ and $+20$ V for both TJ1 and TJ2. One can immediately see that the basic magnetic-field-periodic change in $dI/dV$ is very stable at all $V_{\rm G}$ values. However, upon close inspection, the occasional breaking of exact electron-hole symmetry is visible. Particularly in TJ1, the tunnel-coupling of the electrode to the electron and hole states in the TI surface is modified by back gating: At $V_{\rm G}=-10\,\mathrm{V}$, the electron-like bound states give a stronger signal than the hole-like ones, whereas the situation is reversed for $V_{\rm G}=+15\,\mathrm{V}$. Similar behaviour has been reported for comparable devices, e.g., based on carbon nanotubes[30], and it was explained to be due to state-dependent tunnel couplings. Overall, we conclude that the gap closing is robust with respect to changes of the chemical potential indicating that we observe a topological phase-transition.
@@
In passing, we mention that in a device where the tunnel barrier was not in a favourable condition to probe the TI surface, we observed strongly dispersing bound states in a very narrow range of $V_{\rm G}$ associated with a spurious ZBCP occurring at a particular $V_{\rm G}$ value (see supplement). These features are consistent with the Andreev quantum-dot states predicted in Ref. 37, which might occur in the dielectric. The observations of two different types of ZBCPs, stable and spurious, give confidence that the former is associated with an intrinsic gap closing.
6 Discussions
@@
Our observation of a stable gap closing and reopening in a TI Josephson junction having only a moderate transparency cannot be understood in terms of conventional Andreev bound states and gives the first direct proof of the topological phase-transition predicted by the Fu-Kane theory. To firm up the appearance of topological superconductivity, it would be an interesting future experiment to employ a charge sensing technique[38] in order to detect the parity change in the junction before and after the gap closing. Of course, the definitive proof of topological superconductivity is the observation of Majorana zero-modes, well-protected by a substantial excitation gap, and the demonstration of their non-Abelian exchange statistics through braiding. In a TI Josephson junction, well-protected Majorana zero-modes can in principle be generated in Josephson vortex-cores as we discuss below. Once such protected Majorana zero-modes are generated, they can be physically moved along the junction in a controlled manner[39]; this degree of freedom allows us to conceive braiding operations in a three-slit Josephson junction[39] and the resulting change in the parity can be detected by charge sensing[38]. Another possibility to confirm the non-Abelian statistics is to build a Josephson-vortex interferometer proposed by Grosfeld and Stern[40], which detects the result of braiding using the Aharonov-Casher effect.
@@
valid in the short-junction regime $L\lesssim\hbar v_{\rm F}/\Delta$ ($L$ is the true length of the junction). For large $\mu\gg\Delta$, the formula goes over to the simple expression $v_{\rm M}\simeq v_{\rm F}(\Delta/\mu)^{2}$ given in Ref. 26. The other important quantity is the energy of the first excited state, which is given as $E_{1}\approx\sqrt{2\pi v_{\rm M}\Delta/l_{B}}$ near $\varphi=(2n+1)\pi$ [26]. Note that reducing $\mu$ enhances $v_{\rm M}$, and thus $E_{1}$ and $\lambda_{\rm B}$. However, $\lambda_{\rm B}$ must be shorter than $l_{B}/2$ to avoid the overlap of neighbouring Majorana zero-modes which bounds $\mu$ from below. To see this competition, we theoretically consider the case of using Nb for the SC electrodes. We calculate the dependencies of $\lambda_{\rm B}$ and $E_{1}$ on $\mu$ and $B$. The results, shown in Figs. 2c and 2d, indicate that a moderately large $\mu$ is indeed helpful to realize a short $\lambda_{\rm B}$ and the magnetic field helps to obtain a usefully large $E_{1}$. In Fig. 2d, we draw the boundaries to guarantee a sufficient separation of Majorana zero-modes ($\lambda_{\rm B}/l_{B}<0.2$, blue line) and a sufficiently large excitation gap ($E_{1}/\Delta>0.2$, green line). In the region between the blue and green lines, see Fig. 2d, Majorana zero-modes are well-protected. In fact, these estimates suggest that it is not necessary to tune $\mu$ very close to the Dirac point as long as a magnetic field of the order of 100 mT is used to generate Josephson vortices — an encouraging result for the prospects of Majorana zero-modes in this platform. It is prudent to note that the soft gap, which we observed in our devices, is a problem for braiding, and a hard gap must be realized by improving the interface transparency between the TI and a superconductor. In this regard, we envision the use of the self-epitaxy technique discovered for proximitising (Bi1-xSbx)2Te3 thin films with the PdTe2 superconductor, because a very high transparency of 0.96 was achieved with this technique[41].
@@
In conclusion, we have observed periodic gap-closing and reopening in a phase-biased TI Josephson junction via tunnelling spectroscopy. We have shown that this phenomenon is robust as a function of an extended parameter range in the chemical potential and the perpendicular magnetic field. The robustness of the observed gap closing and reopening is in sharp contrast to the signal expected for a conventional moderate-transparency Josephson junction or other types of zero-energy states which originate from Andreev quantum-dot states[37]. The robustness of the gap closing and reopening is a key feature of the topological phase-transition that occurs in the 1D Majorana state as a function of $\varphi$. In the magnetic-field range where a couple of Josephson vortices nucleate in the junction, we have observed the locality of the gap closing implying a confinement of the zero-energy mode[26]. However, within the magnetic-field range where Al remains superconducting, an abundance of low-lying excited states arises and this leads to spectra that are indistinguishable from the gap closing of an extended 1D Majorana mode. We have discussed the necessary parameter range to realize well-protected Majorana zero-modes in the Josephson vortex core. Intriguingly, we found that maintaining a chemical potential $\mu$ not too close to the Dirac point is crucial to avoid Majorana delocalisation; a $\mu$ of approximately 25 meV facilitates the realization of well-protected Majorana zero-modes in a magnetic field of 100 mT accessible with a Nb electrode. Considering the high controllability of the Josephson vortex, this result is promising for braiding experiments. Overall, our experiment elucidates the topological nature of a TI Josephson junction and underscores the realistic prospect of generating well-protected Majorana zero-modes in this system.
{methods}
Materials and device fabrications: Bulk single crystals of BiSbTeSe2 were synthesized by the modified Bridgman method using high-purity (99.9999%) Bi, Sb, Te, and Se as starting materials by following the recipe described in Ref. 28. We mechanically exfoliated ultra-thin BiSbTeSe2 flakes from bulk single crystals on degenerately-doped Si wafers covered by 280 nm SiO2. The latter acts as dielectric for the back gate. BiSbTeSe2 flakes suitable for device fabrication were identified under an optical microscope. The junctions were defined by electron beam lithography. PMMA A4 resist was exposed using a Raith PIONEER Two system. After development, residual resist in the contact areas was removed by a gentle oxygen plasma treatment. Subsequent cleaning in a hydrochloric acid solution shortly before metallisation ensured the removal of any oxide layer. Then, 3 nm + 45 nm Ti/Al contacts were thermally evaporated after a gentle in-situ Ar cleaning and the device was exposed to air to form a native Al2O3 layer. An additional atomic layer deposition of a thin Al2O3 layer onto the whole device was performed at 80 C using Ultratec Savannah S200 to form a tunnel barrier. For the fabrication of the tunnel electrodes, another electron beam lithography was performed to define the pattern. A 5-nm-thick Pt layer was sputter-deposited to act as wetting layer, followed by sputtering an additional 45-nm-thick Au layer. The contact resistances achieved by this procedure are typically larger than $100\mathrm{k\Omega}$, ensuring electron tunnelling to be the prevalent form of transport. The precise device geometry was determined by scanning electron microscopy after the measurements were finished.
Measurements: Transport measurements were performed at base temperature ($\sim$50 mK) of our dry dilution refrigerator (Oxford Instruments TRITON 300). Electrical lines incorporate RC and copper-powder filters to reduce noise. Using a two-probe configuration, the differential conductance $dI/dV$ was measured between each individual tunnel electrode and the SC electrode with a standard low-frequency lock-in technique using a Basel instruments current preamplifier. A small DC-offset introduced by the preamplifier was numerically corrected for after the measurements. An AC voltage of 5–10 $\mu$V was superimposed to a DC bias voltage. For investigations of the magnetic field dependence, we used our integrated 6/1/1-T superconducting vector magnet.
Simulations: To qualitatively reproduce the conductance spectrum of the TI Josephson junction, a transport simulation was performed using the Python library Kwant[42]. As a model for the three-dimensional strong TI, the Bernevig-Hughes-Zhang model was chosen and implemented on a cubic lattice. Superconductivity was included in the Bogoliubov–de-Gennes formalism. The system has been extended infinitely along the junction and probed by a tunnel lead attached perpendicularly to the junction. For more information see supplement.
{addendum}
This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant Agreement No. 741121) and was also funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - Cluster of Excellence Matter and Light for Quantum Computing (ML4Q) EXC 2004/1 - 390534769, as well as by the DFG under CRC 1238 - 277146847 (Subprojects A04 and B01).
MB grew the BSTS2 crystals, JS and ENN fabricated the devices, performed the experiments and analysed the data together with YA. AM and AZ performed numerical calculations, FH and YA conceived and supervised the project. JS, FH, and YA wrote the manuscript with input from all authors.
The authors declare no competing interests.
Correspondence and requests for materials should be addressed to Y.A. (ando@ph2.uni-koeln.de).
@@
Data availability: The data that support the findings of this study are available at the online depository zenodo with the identifier 10.5281/zenodo.11615514 and in Supplementary Information.

### v2

@@ IV Locality of the gap closing in perpendicular magnetic fields
Next, we study the behavior of our device in higher magnetic fields and compare the LDOS probed at TJ1 and TJ2. When a perpendicular magnetic field $B$ is applied to the junction, $\varphi$ becomes dependent on the position along the junction ($y$ direction). This locality is reflected in the bound-state spectrum, as recently demonstrated in a semiconductor 2DEG-based junction [38]. If $B$ is strong enough to nucleate Josephson vortices, the position where the local phase difference becomes $\varphi=(2n+1)\pi$ is identified as the Josephson vortex-core, and in TI Josephson junctions, the gapless 1D mode in the Fu-Kane theory for global $\varphi=(2n+1)\pi$ turns into a Majorana zero-mode confined to the Josephson vortex-core [27]. The concept of this Majorana zero-mode generation is sketched in Fig. 3a. A spatially varying phase $\varphi$ along a junction caused by $B$ leads to a position-dependent gap function $m\propto\cos\frac{\varphi(y)}{2}$ as indicated by the color code [27]. Associated with the sign change in the gap function, the gap closes locally at $\varphi(y)=(2n+1)\pi$ (white area), corresponding to the core of the Josephson vortex, and a Majorana zero-mode (green dots) nucleates that is confined along the junction by the phase gradient. We note that the topologically-protected nature of the gap closing is intuitively understandable in this situation: The region of the junction with $-\pi<\varphi(y)<\pi$ (red) has the even-parity ground state with $\mathbb{Z}_{2}$ index of 0, whereas the region with $\pi<\varphi(y)<3\pi$ (blue) has the odd-parity ground state with $\mathbb{Z}_{2}$ index of 1; therefore, the boundary between the two regions must become gapless for the topological invariant to change.
Figure 4: Stability of the gap closing and reopening against a change in the chemical potential. a, Differential conductance at a high bias voltage of 1 mV as a function of globally-applied back-gate voltage $V_{\rm G}$. For both tunnel contacts TJ1 and TJ2, the overall conductance is reduced for more negative gate voltages, indicating an $n$-type regime. The vertical dashed lines denote the gate voltages for which the continuous evolution of the spectrum is shown in panels c and d. b, Normalized $dI/dV$, denoted $\left[\frac{dI}{dV}\right]_{\rm norm}$, for TJ1 mapped in the $V_{\rm b}$ vs $V_{\rm G}$ plane for $\varphi=0$ (upper panel) and $\varphi=\pi$ (lower panel). The normalization was done at each $V_{\rm G}$ by using the $dI/dV$ value at $V_{\rm b}=0.3$ mV as reference. c,d, Evolution of the $B$-dependence of the spectrum for a wide range of $V_{\rm G}$ from $-20$ to $+20$ V for TJ1 (c) and TJ2 (d). The $V_{\rm G}$ values selected for the presentation are equally spaced.
@@
We indeed observed the locality of this gap closing. The top and bottom panels of Fig. 3c show the $dI/dV$ data simultaneously measured at TJ1 and TJ2 as a function of $V_{b}$ and $B$ in the 5 – 6 mT range. Both junctions present a comparable evolution of their bound state spectrum, but there is a clear phase shift and their periods are slightly different. For TJ2 we find a periodicity of $200\,\mathrm{\mu T}$, smaller than that of TJ1 ($230\,\mathrm{\mu T}$). This difference in the period is consistent with the sample geometry (see supplement for details) in which the effective junction area considering the flux focusing effect [37, 38, 39] is about $\sim$10of the loop area. To understand the LDOS at TJ1 and TJ2, we calculated the expected texture of the gap function based on the device geometry. In our junctions, the self-inductance is negligible and its width is much shorter than the Josephson penetration depth (see supplement), justifying the description of the phase gradient in terms of
@@
with $A$ the loop area, $\Phi_{0}=\frac{h}{2e}$ the magnetic flux quantum, and $L_{\rm eff}$ the effective length of the junction taking the flux-focusing effect into account. The evolution of the gap function $m\propto\cos\frac{\varphi(y)}{2}$ is calculated from this formula with the loop and junction area as fitting parameters (see supplement) and visualized in the middle panel of Fig. 3c. We show snapshots of the phase-texture in the junction for a range of magnetic fields corresponding to the experimental situation. In this range, there are two to three Josephson vortices inside the junction, whose position is tuned by the global perpendicular magnetic field. They enter the junction from the bottom where TJ2 is located, and move upwards to TJ1 with increasing magnetic field. Whenever a Josephson vortex-core crosses the location of a tunnel junction, a gap closing is observed. This demonstrates the local nature of the gap closing in the presence of Josephson vortices. Note that in a similar experiment on a non-topological Josephson junction [38], the gap was found to remain open at the Josephson vortex-core. Our data shown in Fig. 3c was taken at an applied back-gate voltage of $V_{\rm G}=-5$ V, and we chose the magnetic field range of 5 – 6 mT for the visibility of the phase gradient. The same pattern was found to continue into higher magnetic fields until the Al superconductivity was suppressed at 15 mT (see supplement).
@@
We note that the spectra shown in Fig. 3c essentially look the same as those in Fig. 1e. This is because the Majorana zero-mode in the Josephson vortex-core is not well-separated in energy. For the conditions realized, it is accompanied by a quasi-continuum of low-energy bound states [27]. Taking the smearing due to the finite-temperature into account, the local spectrum becomes a broad peak, indistinguishable from the Fu-Kane situation. In Fig. 3b, we show the magnitude of the gap function near a Josephson vortex-core and the finite-energy bound states calculated from the model in Ref. 27, along with the spatial extent of the expected Majorana zero-mode. For the parameters relevant to our device with $\mu$ = 25 meV, the calculated finite-energy bound states form a quasi-continuum with a minigap of only 2% of $\Delta$ (i.e. 3 $\mu$eV), which is of the order of the thermal energy at the base temperature.
V Stability of the gap closing
@@
In the next step, we check the robustness of the observed features with respect to the chemical potential, as tuned by a global back gate. Figure 4a shows the $dI/dV$ data in $B$ = 0 T at a high bias voltage ($V_{\rm b}$ = 1 mV) as a function of the back-gate voltage $V_{G}$ for both TJ1 and TJ2. The overall tendency is that the conductance decreases for more negative gate-voltages, indicating that the TSS is in the $n$-type regime. Superimposed to this general trend, we observe broad peaks (typical peak width of a few V), which are commonly found in dielectric-based tunnelling devices and conventionally explained by resonant states in the dielectric [31]. We can use their dispersion to estimate the relation between $V_{G}$ and $\mu$ (see supplement) and find $\Delta\mu/\Delta V_{G}\approx 1-2\,\mathrm{meV/V}$, consistent with results reported in our earlier work [37]. This implies that the chemical potential $\mu$ changes by at least $40\,\mathrm{meV}$ in the gating range studied here.
Figure 4b presents the $V_{\rm G}$-dependencies of the normalized conductance spectra obtained at TJ1. Here, we show the evolutions of the maximum excitation gap for $\varphi=0$ (upper panel) and the spectrum observed for $\varphi=\pi$ (lower panel) in a wide range of $V_{\rm G}$ from $-20$ to $+20$ V. The maximum excitation gap is very stable, indicating that the resonance states due to impurities are not affecting the proximity-induced gap. The gap closing at $\varphi=\pi$ is also very stable and is observed at all $V_{\rm G}$ values. Note that the occurrence of the broad ZBCP and the details of the LDOS near zero-bias appear to depend on $V_{\rm G}$, but as we noted in Sec. III, these features are expected to be sensitive to the details of the experimental conditions; for example, the $dI/dV$ peak is slightly shifted to the negative $V_{\rm b}$ side at $V_{\rm G}\approx$ 12 – 20 V, which is probably due to an influence of the strong resonance state located around 13 V. What is important in these data is the robustness of the gap closing, which is observed for the whole range of $V_{\rm G}$.
@@
In Figs. 4c and 4d, we show the $B$-dependencies of the normalized conductance spectra for equally-spaced $V_{\rm G}$ values between $-20$ and $+20$ V for both TJ1 and TJ2. One can immediately see that the basic magnetic-field-periodic change in $dI/dV$ is very stable at all $V_{\rm G}$ values. Even though occasional breaking of exact electron-hole symmetry is visible upon close inspection, similar behavior has been reported for comparable devices, e.g., based on carbon nanotubes [33], and it was explained to be due to state-dependent tunnel couplings. Overall, we conclude that the gap closing is robust with respect to changes of the chemical potential, consistent with the TSS origin.
VI Discussions
@@
It is prudent to mention that a gap closing at $\varphi=\pi$ can occur in some non-topological scenarios. A quantum-dot state, symmetrically coupled to SC electrodes and tuned to resonance, leads to a zero-energy state [40]. This highly fine-tuned case was in fact observed by us in a device where the tunnel barrier was not in a favourable condition to probe the TSS. We found strongly dispersing bound states in a very narrow range of $V_{\rm G}$ associated with a spurious ZBCP occurring at a particular $V_{\rm G}$ value (see supplement). These features are consistent with the Andreev quantum-dot states [40] which might occur in the dielectric. The robust gap closing discussed in the previous section is clearly of different origin. Nonetheless, the sharp Andreev quantum-dot states are useful for estimating the energy resolution of our setup to be about 30 $\mu$eV (see supplement). The spurious ZBCP in this case also tells us that one should not look for a ZBCP but should focus on the gap closing to elucidate the robustness of the phenomenon.
Figure 5: Stability of the Majorana zero-modes in the putative case with Nb electrodes. a,b, Calculated dependencies of the Majorana bound state width $\lambda_{\rm B}$ and the first excited state energy $E_{1}$ on the chemical potential $\mu$ and the perpendicular magnetic field $B$ for the Majorana zero-mode generated at the Josephson vortex-core in a putative device with Nb as the SC electrodes, which will sustain high magnetic fields. We used $L$ = 50 nm, $v_{\rm F}$ = 5.5 $\times$ 105 m/s, and $\Delta$ = 1 meV. The blue and green lines in panel b are the boundaries to guarantee a sufficient separation of Majorana zero-modes ($\lambda_{\rm B}/l_{B}<0.2$ obtained from panel a) and a sufficiently large excitation gap ($E_{1}/\Delta>0.2$), respectively. Majorana zero-modes are well-protected in the relatively broad region between the blue and green lines.
In the literature, the gap closing at $\varphi=\pi$ has been reported for diffusive long SNS junctions [41, 32]. In such a system, disorder localizes most channels, but it also leads to the appearance of a few high-transmission channels [42], which can be responsible for a gap closing. In the case of our junctions, the mean-free path $l_{e}\approx$ 40 nm has been estimated for bulk-insulating Bi1.5Sb0.5Te1.7Se1.3 crystals [43], which is a bit shorter but comparable to our junction length (50 nm), putting our device in a quasi-ballistic regime. (Remember that our device is in the short-junction limit where the Fu-Kane theory is applicable.) It is possible that some nearly-longitudinal Andreev modes with large $k_{y}$ traveling for a long distance present the gap closing phenomenology of the diffusive scenario. Nevertheless, such a gap closing occurs on top of the intrinsic gap closing guaranteed by the spin-momentum locking for the transverse mode [26]. Hence, the consideration of the diffusive contribution does not change the conclusion that the gap closing predicted by Fu and Kane is being observed in our device, unless the junction is not coupled to the TSS (which can be dismissed based on Ref. 37). To better address the question of which states contribute to what extent to the gap-closing, employing high-frequency measurements to achieve a higher resolution [44] would be useful. Furthermore, employing a charge-sensing technique [45] to detect the parity change in the junction before and after the gap closing would unambiguously prove the topological phase transition.
@@
A definitive proof of Majorana zero-modes is the demonstration of their non-Abelian exchange statistics through braiding. In a TI Josephson junction, well-protected Majorana zero-modes can in principle be generated in Josephson vortex-cores as we discuss below. Once such protected Majorana zero-modes are generated, they can be physically moved along the junction in a controlled manner [46]; this degree of freedom allows us to conceive braiding operations in a three-slit Josephson junction [46] and the resulting change in the parity can be detected by charge sensing [45]. Another possibility to confirm the non-Abelian statistics is to build a Josephson-vortex interferometer proposed by Grosfeld and Stern [47], which detects the result of braiding using the Aharonov-Casher effect.
@@
valid in the short-junction regime $L\lesssim\hbar v_{\rm F}/\Delta$ ($L$ is the true length of the junction). For large $\mu\gg\Delta$, the formula goes over to the simple expression $v_{\rm M}\simeq v_{\rm F}(\Delta/\mu)^{2}$ given in Ref. 27. The other important quantity is the energy of the first excited state, which is given as $E_{1}\approx\sqrt{2\pi v_{\rm M}\Delta/l_{B}}$ near $\varphi=(2n+1)\pi$ [27]. Note that reducing $\mu$ enhances $v_{\rm M}$, and thus $E_{1}$ and $\lambda_{\rm B}$. However, $\lambda_{\rm B}$ must be shorter than $l_{B}/2$ to avoid the overlap of neighbouring Majorana zero-modes which bounds $\mu$ from below. To see this competition, we theoretically consider the case of using Nb for the SC electrodes. We calculate the dependencies of $\lambda_{\rm B}$ and $E_{1}$ on $\mu$ and $B$. The results, shown in Figs. 5a and 5b, indicate that a moderately large $\mu$ is indeed helpful to realize a short $\lambda_{\rm B}$ and the magnetic field helps to obtain a usefully large $E_{1}$. In Fig. 5b, we draw the boundaries to guarantee a sufficient separation of Majorana zero-modes ($\lambda_{\rm B}/l_{B}<0.2$, blue line) and a sufficiently large excitation gap ($E_{1}/\Delta>0.2$, green line). In the region between the blue and green lines, see Fig. 5b, Majorana zero-modes are well-protected. In fact, these estimates suggest that it is not necessary to tune $\mu$ very close to the Dirac point as long as a magnetic field of the order of 100 mT is used to generate Josephson vortices — an encouraging result for this platform. It is prudent to note that the soft gap, which we observed in our devices, is a problem for braiding, and a hard gap must be realized by improving the interface transparency between the TI and a superconductor.
@@
In conclusion, we have observed periodic gap-closings and reopenings in a phase-biased TI Josephson junction, which is consistent with the existence of perfectly-transmitted Andreev modes guaranteed by the spin-momentum locking of the TSS. This phenomenon is robust as a function of an extended parameter range in the chemical potential and the perpendicular magnetic field, which support its TSS origin. The spin-non-degeneracy of the TSS makes this gap-closing and reopening to be topological phase transition accompanied by parity change. In the magnetic-field range where a couple of Josephson vortices nucleate in the junction, we observed the locality of the gap closing implying a confinement of the zero-energy mode [27]. However, within the magnetic-field range where Al remains superconducting, an abundance of low-lying excited states arises and this leads to spectra that are indistinguishable from the gap closing near zero field. We have discussed the necessary parameter range to realize well-protected Majorana zero-modes in the Josephson vortex core. Intriguingly, we found that maintaining a chemical potential $\mu$ not too close to the Dirac point is crucial for avoiding Majorana delocalisation; a $\mu$ of approximately 25 meV facilitates the realization of well-protected Majorana zero-modes in a magnetic field of 100 mT accessible with a Nb electrode. Considering the high controllability of the Josephson vortex, this result is promising for braiding experiments. Overall, our experiment demonstrates a highly promising nature of the TI Josephson junctions for the Majorana research thanks to the naturally-existing spin-momentum locking of the TSS.
@@
Data availability: The data and codes that support the findings of this study are available at the online depository zenodo with the identifier 10.5281/zenodo.11615512 and Supplemental Material.
Acknowledgements.
We thank Liang Fu, Max Geier, Michal Papaj, and Henry Legg for fruitful discussions. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant Agreement No. 741121) and was also funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - Cluster of Excellence Matter and Light for Quantum Computing (ML4Q) EXC 2004/1 - 390534769, as well as by the DFG under CRC 1238 - 277146847 (Subprojects A04 and B01).